# Distribution estimation via Flow Matching with Lipschitz guarantees

### v1

@@ 7.2 Proofs of Section˜4
$\big{(}1+\frac{\sigma_{t}^{2}}{t^{2}}\big{)}^{2}\mathbb{E}_{p_{Z}}\Big{[}\Big{(}Z_{i}-\mathbb{E}[Z_{i}]\Big{)}^{2}\Big{]}$ $=\big{(}1+(\frac{\sigma_{t}}{t})^{2}\big{)}+(\frac{\sigma_{t}}{t})^{2}\mathbb{E}_{p_{Z}}[\partial^{2}_{z_{i}}a(\frac{x-\sigma_{t}z}{t})]$
$-\big{(}\frac{\sigma_{t}}{t}\big{)}^{2}\mathbb{E}_{p_{Z}}\Big{[}\Big{(}\partial_{z_{i}}a(\frac{x-\sigma_{t}Z}{t})\Big{)}^{2}\Big{]}-\big{(}1+\frac{\sigma_{t}^{2}}{t^{2}}\big{)}^{2}\Big{(}\mathbb{E}[Z_{i}]-x_{i}\frac{\sigma_{t}}{t^{2}(1+(\frac{\sigma_{t}}{t})^{2})}\Big{)}^{2}$
$\quad-2\big{(}1+\frac{\sigma_{t}^{2}}{t^{2}}\big{)}\frac{\sigma_{t}}{t}\mathbb{E}_{p_{Z}}\Big{[}(Z_{i}-\mathbb{E}[Z_{i}])\partial_{z_{i}}a(\frac{x-\sigma_{t}Z}{t})\Big{]}$
$\quad-2\big{(}1+\frac{\sigma_{t}^{2}}{t^{2}}\big{)}\frac{\sigma_{t}}{t}(\mathbb{E}[Z_{i}]-x_{i}\frac{\sigma_{t}}{t^{2}(1+(\frac{\sigma_{t}}{t})^{2})})\mathbb{E}_{p_{Z}}\Big{[}\partial_{z_{i}}a(\frac{x-\sigma_{t}Z}{t})\Big{]}.$
@@
$\big{(}1+\frac{\sigma_{t}^{2}}{t^{2}}\big{)}^{2}\mathbb{E}_{p_{Z}}\Big{[}\Big{(}Z_{i}-\mathbb{E}[Z_{i}]\Big{)}^{2}\Big{]}$ $\leq\big{(}1+(\frac{\sigma_{t}}{t})^{2}\big{)}+(\frac{\sigma_{t}}{t})^{2}L+2\big{(}1+\frac{\sigma_{t}^{2}}{t^{2}}\big{)}\frac{\sigma_{t}}{t}L\sqrt{\mathbb{E}_{p_{Z}}\big{[}(Z_{i}-\mathbb{E}[Z_{i}])^{2}\big{]}}$
$\quad+2\big{(}1+\frac{\sigma_{t}^{2}}{t^{2}}\big{)}\frac{\sigma_{t}^{2}}{t^{2}}\frac{L^{2}}{\big{(}1+\frac{\sigma_{t}^{2}}{t^{2}}\big{)}},$
@@
$\big{(}1+\frac{\sigma_{t}^{2}}{t^{2}}\big{)}^{2}\mathbb{E}_{p_{Z}}\Big{[}\Big{(}Z_{i}-\mathbb{E}[Z_{i}]\Big{)}^{2}\Big{]}$ $\geq\big{(}1+(\frac{\sigma_{t}}{t})^{2}\big{)}-(\frac{\sigma_{t}}{t})^{2}L-\big{(}\frac{\sigma_{t}}{t}\big{)}^{2}L^{2}-\big{(}1+\frac{\sigma_{t}^{2}}{t^{2}}\big{)}^{2}\Big{(}\frac{\sigma_{t}}{t}\frac{L}{\big{(}1+\frac{\sigma_{t}^{2}}{t^{2}}\big{)}}\Big{)}^{2}$ (26)
@@
With the same reasoning via the Poincaré constant of a Gaussian and the Holley-Stroock perturbation principle, we conclude that
$\mathbb{E}_{p_{Z}}\big{[}(Z_{i}-\mathbb{E}[Z_{i}])^{2}\big{]}\leq e^{4L}.$
As $\big{(}1+\frac{\sigma_{t}^{2}}{t^{2}}\big{)}^{2}>0$ and
$\frac{1}{1+\frac{\sigma_{t}^{2}}{t^{2}}}=1-\frac{\frac{\sigma_{t}^{2}}{t^{2}}}{1+\frac{\sigma_{t}^{2}}{t^{2}}},$
dividing the first term in the upper and lower bound loosens the bound further. Thus we obtain
$\mathbb{E}_{p_{Z}}\big{[}\big{(}Z_{i}-\mathbb{E}[Z_{i}]\big{)}^{2}\big{]}=1+O(\frac{\sigma_{t}}{t}).$
Multiplying with $\frac{\sigma_{t}^{2}}{t}$ yields the result for the variances via (21):
$\operatorname{Var}(Y^{x,t}_{i})=\big{(}\frac{\sigma_{t}}{t}\big{)}^{2}\big{(}1+O\big{(}\frac{\sigma_{t}}{t}\big{)}).$
Property (I):
For the covariances, we need to find a stricter upper bound. We are going to use (Menz, 2014, Theorem 2.3) applied to the component functions $f_{i}\colon\mathbb{R}^{d}\rightarrow\mathbb{R},\;f_{i}(y)=y_{i}$. First we verify that the assumptions are fulfilled in our setting. By construction, the distribution of $Y_{i}^{x,t}$ is a bounded perturbation of a Gaussian. Hence Assumption 2.2 in Menz (2014) is satisfied.
@@
where $a(\cdot|y_{1},...,y_{i-1},y_{i+1},...,y_{d})\coloneqq a(y_{1},...,y_{i-1},\cdot,y_{i+1},...,y_{d})$ and $i\in\{1,...,d\}$, we can use the same arguments as in the proof of property (III) to obtain
$\rho_{i}^{t}\geq e^{-4L}(1+\frac{t^{2}}{\sigma_{t}^{2}}).$
Note that $a(\cdot|y_{1},...,y_{i-1},y_{i+1},...,y_{d})$ is still bounded. Furthermore, we know that the off-diagonal entries of the Hessian of the $\log$-density are bounded by $L$. To profit from easier notation later, we define the matrix
@@
$e^{-4L}(1+\frac{t^{2}}{\sigma_{t}^{2}})>L(d-1)\quad\Longleftarrow\quad\frac{\sigma_{t}}{t}<\frac{1}{\sqrt{e^{4L}L(d-1)-1}}.$ (28)
@@
We conclude that
$|\operatorname{Cov}(Y^{x,t})_{ij}|\leq((e^{-4L}\frac{t^{2}}{\sigma_{t}^{2}}I_{d}+A_{t})^{-1})_{ij}.$
Now
@@
As $\rho_{i}^{t}-e^{-4L}\frac{t^{2}}{\sigma_{t}^{2}}=e^{-4L}$ we conclude that the largest eigenvalue is bounded by
$\lambda_{\max}(A_{t})\leq e^{-4L}+L(d+1).$
We therefore deduce that
@@
Using the convergence of the geometric series, we get that
$\sum_{k=2}^{\infty}\big{(}\frac{\sigma_{t}}{t}\big{)}^{k}=\frac{\frac{\sigma_{t}^{2}}{t^{2}}}{1-\frac{\sigma_{t}}{t}}\leq\frac{\sigma_{t}^{2}}{t^{2}}.$
Inserting everything into (29), we obtain
@@
If we choose $t^{*}\geq\frac{1}{2}$, we know that $\frac{\sigma_{t}^{2}}{t^{2}}\leq 2\sigma_{t}^{2}$. This gives the smaller bound on the covariances.
7.3 Proofs of Section˜5
Proof of Theorem˜5.2.
@@ 7.3 Proofs of Section˜5
Then we use Lemma˜2.1,
$\mathbb{E}_{\begin{subarray}{c}t\sim\mathcal{U}[0,1]\\ X_{t}\sim p_{t}\end{subarray}}\big{[}|v_{t}(X_{t})-\hat{v}_{t}(X_{t})|^{2}\big{]}=\mathbb{E}_{\begin{subarray}{c}t\sim\mathcal{U}[0,1],\\ Y\sim p^{*},\\ X_{t}\sim p_{t}\left(\cdot|Y\right)\end{subarray}}\big{[}\left|\hat{v}_{t}(X_{t})-v_{t}\left(X_{t}|Y\right)\right|^{2}\big{]}-C,$ (32)
where
$C\coloneqq\mathbb{E}_{\begin{subarray}{c}t\sim\mathcal{U}[0,1],\\ Y\sim p^{*},\\ X_{t}\sim p_{t}\left(\cdot|Y\right)\end{subarray}}[|v_{t}(X_{t})-v_{t}(X_{t}|Y)|^{2}].$
@@
$\mathbb{E}_{\begin{subarray}{c}t\sim\mathcal{U}[0,1],\\ Y\sim p^{*},\\ X_{t}\sim p_{t}\left(\cdot|Y\right)\end{subarray}}\big{[}\left|\hat{v}_{t}(X_{t})-v_{t}\left(X_{t}|Y\right)\right|^{2}\big{]}$ $=\int_{A}\int_{0}^{1}\int|\hat{v}_{t}(x)-v_{t}(x|y)|^{2}p_{t}(x|y)\;\mathrm{d}x\;\mathrm{d}t\;p^{*}(y)\;\mathrm{d}y$
$\quad+\int_{\mathbb{R}^{d}\setminus A}\int_{0}^{1}\int|\hat{v}_{t}(x)-v_{t}(x|y)|^{2}p_{t}(x|y)\;\mathrm{d}x\;\mathrm{d}t\;p^{*}(y)\;\mathrm{d}y.$
Note that
$\int_{A}\int_{0}^{1}\int|\hat{v}_{t}(x)-v_{t}(x|y)|^{2}p_{t}(x|y)\;\mathrm{d}x\;\mathrm{d}t\;p^{*}(x)\;\mathrm{d}y=\mathbb{E}_{Y\sim\mathbb{P}^{*}}[g(\hat{v},Y)\mathds{1}_{Y\in A}].$
@@
$|v_{t}^{j}(x)|$ $=\Big{|}\int v^{j}_{t}(x|y)\frac{p_{t}(x|y)}{\int p_{t}(x|z)p^{*}(z)\;\mathrm{d}z}p^{*}(y)\;\mathrm{d}y\Big{|}$
$=\Big{|}\int\big{(}\frac{\sigma^{\prime}_{t}}{\sigma_{t}}x_{j}+(1-\frac{\sigma^{\prime}_{t}}{\sigma_{t}}t)y_{j}\big{)}\frac{p_{t}(x|y)}{\int p_{t}(x|z)p^{*}(z)\;\mathrm{d}z}p^{*}(y)\;\mathrm{d}y\Big{|}$
@@
$\leq\big{|}\frac{\sigma^{\prime}_{t}}{\sigma_{t}}\big{|}|x_{j}|+e^{2L}\big{|}\big{(}1-\frac{\sigma^{\prime}_{t}}{\sigma_{t}}t\big{)}\big{|}(\frac{t}{t^{2}+\sigma_{t}^{2}}|x_{j}|+(1+\frac{t^{2}}{\sigma_{t}^{2}})^{-\frac{1}{2}})$
$=\big{(}\big{|}\frac{\sigma^{\prime}_{t}}{\sigma_{t}}\big{|}+e^{2L}\big{|}\big{(}1-\frac{\sigma^{\prime}_{t}}{\sigma_{t}}t\big{)}\big{|}\frac{t}{t^{2}+\sigma_{t}^{2}}\big{)}|x_{j}|+\big{|}\big{(}1-\frac{\sigma^{\prime}_{t}}{\sigma_{t}}t\big{)}\big{|}(1+\frac{t^{2}}{\sigma_{t}^{2}})^{-\frac{1}{2}}.$
As
$\max_{t\in[0,1]}\big{|}\frac{\sigma^{\prime}_{t}}{\sigma_{t}}\big{|}=\log(\sigma_{\min}^{-1}),\quad(1+\frac{t^{2}}{\sigma_{t}^{2}})^{-\frac{1}{2}}\leq 1$
and due to Lemma˜7.10
$\frac{t}{t^{2}+\sigma_{t}^{2}}\leq\max(\log(\sigma_{\min}^{-1}),e^{2})$
we obtain for the maximum over $t\in[0,1]$ for $n$ big enough
$|v_{t}^{j}(x)|$ $\lesssim\log(n)^{2}|x_{j}|+\log(n).$
The constants lead to the bound in (15). An analogous calculation shows that, for $t\in[0,1]$, the norm of $v_{t}$ is bounded by
@@
$g(v,y)$ $=\int_{0}^{1}\int|v_{t}(x)-v_{t}(x|y)|^{2}p_{t}(x|y)\;\mathrm{d}x\;\mathrm{d}t$
$\lesssim\log(n)^{6}+\int_{0}^{1}\int|v(x|y)|^{2}p_{t}(x|y)\;\mathrm{d}x\;\mathrm{d}t.$
@@
$\int|v(x|y)|^{2}p_{t}(x|y)\;\mathrm{d}x$ $=\int\big{|}\frac{\sigma^{\prime}_{t}}{\sigma_{t}}x+\big{(}1-\frac{\sigma^{\prime}_{t}}{\sigma_{t}}t\big{)}y\big{|}^{2}p_{t}(x|y)\;\mathrm{d}x$
$=\big{|}\frac{\sigma^{\prime}_{t}}{\sigma_{t}}\big{|}^{2}\int|x|^{2}p_{t}(x|y)\;\mathrm{d}x+\big{(}1-\frac{\sigma^{\prime}_{t}}{\sigma_{t}}t\big{)}^{2}|y|^{2},$
and as $p_{t}(\cdot|y)$ is the density of a Gaussian with mean $y$ and variance $\sigma_{t}^{2}I_{d}$,
$\int|x|^{2}p_{t}(x|y)\;\mathrm{d}x$ $=t^{2}|y|^{2}+\sigma_{t}^{2}d,$ (34)
we obtain for $g(v,y)$
$g(v,y)$ $\lesssim\log(n)^{6}+\int_{0}^{1}\big{|}\frac{\sigma^{\prime}_{t}}{\sigma_{t}}\big{|}^{2}(t^{2}|y|^{2}+\sigma_{t}^{2}d)\;\mathrm{d}t+|y|^{2}\int_{0}^{1}(1-\frac{\sigma^{\prime}_{t}}{\sigma_{t}}t)^{2}\;\mathrm{d}t\lesssim\log(n)^{6}.$

### v2

@@ 6.1 Proofs of Section˜3
$\Big(1+\frac{\sigma_{t}^{2}}{t^{2}}\Big)^{2}\mathbb{E}_{p_{Z}}\Big[\Big(Z_{i}-\mathbb{E}[Z_{i}]\Big)^{2}\Big]=\Big(1+\Big(\frac{\sigma_{t}}{t}\Big)^{2}\Big)+\Big(\frac{\sigma_{t}}{t}\Big)^{2}\mathbb{E}_{p_{Z}}\Big[\partial^{2}_{z_{i}}a\Big(\frac{x-\sigma_{t}z}{t}\Big)\Big]-\Big(\frac{\sigma_{t}}{t}\Big)^{2}\mathbb{E}_{p_{Z}}\Big[\Big(\partial_{z_{i}}a(\frac{x-\sigma_{t}Z}{t})\Big)^{2}\Big]$
$\quad-\Big(1+\frac{\sigma_{t}^{2}}{t^{2}}\Big)^{2}\Big(\mathbb{E}[Z_{i}]-x_{i}\frac{\sigma_{t}}{t^{2}(1+(\frac{\sigma_{t}}{t})^{2})}\Big)^{2}-2\Big(1+\frac{\sigma_{t}^{2}}{t^{2}}\Big)\frac{\sigma_{t}}{t}\mathbb{E}_{p_{Z}}\Big[(Z_{i}-\mathbb{E}[Z_{i}])\partial_{z_{i}}a\Big(\frac{x-\sigma_{t}Z}{t}\Big)\Big]$
$\quad-2\Big(1+\frac{\sigma_{t}^{2}}{t^{2}}\Big)\frac{\sigma_{t}}{t}\Big(\mathbb{E}[Z_{i}]-x_{i}\frac{\sigma_{t}}{t^{2}(1+(\frac{\sigma_{t}}{t})^{2})}\Big)\mathbb{E}_{p_{Z}}\Big[\partial_{z_{i}}a\Big(\frac{x-\sigma_{t}Z}{t}\Big)\Big].$
@@
$\Big(1+\frac{\sigma_{t}^{2}}{t^{2}}\Big)^{2}\mathbb{E}_{p_{Z}}\big[\big(Z_{i}-\mathbb{E}[Z_{i}]\big)^{2}\big]\leq\Big(1+\Big(\frac{\sigma_{t}}{t}\Big)^{2}\Big)+\Big(\frac{\sigma_{t}}{t}\Big)^{2}L+2\Big(1+\frac{\sigma_{t}^{2}}{t^{2}}\Big)\frac{\sigma_{t}}{t}L\sqrt{\mathbb{E}_{p_{Z}}\big[(Z_{i}-\mathbb{E}[Z_{i}])^{2}\big]}+2\Big(1+\frac{\sigma_{t}^{2}}{t^{2}}\Big)\frac{\sigma_{t}^{2}}{t^{2}}\frac{L^{2}}{\big(1+\frac{\sigma_{t}^{2}}{t^{2}}\big)},$
@@
$\Big(1+\frac{\sigma_{t}^{2}}{t^{2}}\Big)^{2}\mathbb{E}_{p_{Z}}\Big[\Big(Z_{i}-\mathbb{E}[Z_{i}]\Big)^{2}\Big]$ $\geq\Big(1+\Big(\frac{\sigma_{t}}{t}\Big)^{2}\Big)-(\frac{\sigma_{t}}{t})^{2}L-\Big(\frac{\sigma_{t}}{t}\Big)^{2}L^{2}-\Big(1+\frac{\sigma_{t}^{2}}{t^{2}}\Big)^{2}\Big(\frac{\sigma_{t}}{t}\frac{L}{\Big(1+\frac{\sigma_{t}^{2}}{t^{2}}\Big)}\Big)^{2}$ (27)
@@
With the same reasoning via the Poincaré constant of a Gaussian and the Holley-Stroock perturbation principle, we conclude that $\mathbb{E}_{p_{Z}}\big[(Z_{i}-\mathbb{E}[Z_{i}])^{2}\big]\leq e^{4L}.$ As $\big(1+\frac{\sigma_{t}^{2}}{t^{2}}\big)^{2}>0$ and $\frac{1}{1+\frac{\sigma_{t}^{2}}{t^{2}}}=1-\frac{\frac{\sigma_{t}^{2}}{t^{2}}}{1+\frac{\sigma_{t}^{2}}{t^{2}}},$ dividing the first term in the upper and lower bound loosens the bound further. Thus we obtain $\mathbb{E}_{p_{Z}}\big[\big(Z_{i}-\mathbb{E}[Z_{i}]\big)^{2}\big]=1+O(\frac{\sigma_{t}}{t}).$ Multiplying with $\frac{\sigma_{t}^{2}}{t}$ yields the result for the variances via (22) $\operatorname{Var}(Y^{x,t}_{i})=\big(\frac{\sigma_{t}}{t}\big)^{2}\big(1+O\big(\frac{\sigma_{t}}{t}\big)).$
Property (I), $\mathbb{P}^{*}$ of form (13):
For the covariances, we need to find a stricter upper bound. We are going to use Menz (2014, Theorem 2.3) applied to the component functions $f_{i}\colon\mathbb{R}^{d}\rightarrow\mathbb{R},\;f_{i}(y)=y_{i}$. First we verify that the assumptions are fulfilled in our setting. By construction, the distribution of $Y_{i}^{x,t}$ is a bounded perturbation of a Gaussian. Hence Assumption 2.2 in Menz (2014) is satisfied.
@@
where $a(\cdot|y_{1},...,y_{i-1},y_{i+1},...,y_{d})\coloneqq a(y_{1},...,y_{i-1},\cdot,y_{i+1},...,y_{d})$ and $i\in\{1,...,d\}$, we can use the same arguments as in the proof of property (III) to obtain $\rho_{i}^{t}\geq e^{-4L}(1+\frac{t^{2}}{\sigma_{t}^{2}}).$ Note that $a(\cdot|y_{1},...,y_{i-1},y_{i+1},...,y_{d})$ is still bounded. Furthermore, we know that the off-diagonal entries of the Hessian of the $\log$-density are bounded by $L$. To profit from easier notation later, we define the matrix
@@
$e^{-4L}\Big(1+\frac{t^{2}}{\sigma_{t}^{2}}\Big)>L(d-1)\quad\Longleftarrow\quad\frac{\sigma_{t}}{t}<\frac{1}{\sqrt{e^{4L}L(d-1)-1}}.$ (29)
@@
We conclude that $|\operatorname{Cov}(Y^{x,t})_{ij}|\leq((e^{-4L}\frac{t^{2}}{\sigma_{t}^{2}}I_{d}+A_{t})^{-1})_{ij}.$ Now
@@
As $\rho_{i}^{t}-e^{-4L}\frac{t^{2}}{\sigma_{t}^{2}}=e^{-4L}$ we conclude that the largest eigenvalue is bounded by $\lambda_{\max}(A_{t})\leq e^{-4L}+L(d+1).$ We therefore deduce that
@@
Using the convergence of the geometric series, we get that $\sum_{k=2}^{\infty}\big(\frac{\sigma_{t}}{t}\big)^{k}=\frac{\frac{\sigma_{t}^{2}}{t^{2}}}{1-\frac{\sigma_{t}}{t}}\leq\frac{\sigma_{t}^{2}}{t^{2}}.$ Inserting everything into (30), we obtain
@@
If we choose $t^{*}\geq\frac{1}{2}$, we know that $\frac{\sigma_{t}^{2}}{t^{2}}\leq 2\sigma_{t}^{2}$. This gives the smaller bound on the covariances. ∎
Proof of Example˜3.7.
We start combining (4), (10) and the choice of $\mathbb{P}^{*}$ to obtain
$v_{t}(x)=\frac{A(x)}{B(x)},\quad\text{where}\quad A(x)\coloneqq\int_{0}^{1}\Big(\frac{\sigma_{t}^{\prime}}{\sigma_{t}}(x-ty)+y\Big)\exp\Big(-\frac{(x-ty)^{2}}{2\sigma_{t}^{2}}\Big)\;\mathrm{d}y,\quad B(x)=\int_{0}^{1}\exp\Big(-\frac{(x-ty)^{2}}{2\sigma_{t}^{2}}\Big)\;\mathrm{d}y.$
By the quotient rule, we obtain $\partial_{x}v_{t}(x)=\frac{\partial_{x}A(x)}{B(x)}-\frac{\partial_{x}B(x)}{B(x)}\frac{A(x)}{B(x)}.$ Differentiating leads to
$\partial_{x}A(x)$ $=\int_{0}^{1}\frac{\sigma_{t}^{\prime}}{\sigma_{t}}\exp\Big(-\frac{(x-ty)^{2}}{2\sigma_{t}^{2}}\Big)\;\mathrm{d}y-\int_{0}^{1}\Big(\frac{\sigma_{t}^{\prime}}{\sigma_{t}}(x-ty)+y\Big)\frac{x-ty}{\sigma_{t}^{2}}\exp\Big(-\frac{(x-ty)^{2}}{2\sigma_{t}^{2}}\Big)\;\mathrm{d}y,$
$\partial_{x}B(x)$ $=-\int_{0}^{1}\frac{x-ty}{\sigma_{t}^{2}}\exp\Big(-\frac{(x-ty)^{2}}{2\sigma_{t}^{2}}\Big)\;\mathrm{d}y.$
We note that the integrands on the right side are finite due to the exponential decay, which makes the exchange of integration and differentiation valid. Inserting the partial derivatives and using $Y^{x}\sim p_{t}(x|\cdot)p^{*}(\cdot)$
$\partial_{x}v_{t}(x)=\frac{\sigma_{t}^{\prime}}{\sigma_{t}}-\mathbb{E}\Big[\big(\frac{\sigma_{t}^{\prime}}{\sigma_{t}}(x-tY^{x})+Y^{x}\big)\frac{x-tY^{x}}{\sigma_{t}^{2}}\Big]+\mathbb{E}\Big[\frac{x-tY^{x}}{\sigma_{t}^{2}}\Big]\mathbb{E}\Big[\Big(\frac{\sigma_{t}^{\prime}}{\sigma_{t}}(x-tY^{x})+Y^{x}\Big)\Big].$
Rewriting the second term as the covariance and using linearity and the irrelevance constants of the covariance leads to
$\operatorname{Cov}\Big(\Big(\frac{\sigma_{t}^{\prime}}{\sigma_{t}}(x-tY^{x})+Y^{x}\Big),\frac{x-tY^{x}}{\sigma_{t}^{2}}\Big)=\Big(\frac{\sigma_{t}^{\prime}}{\sigma_{t}^{3}}t^{2}-\frac{t}{\sigma_{t}^{2}}\Big)\operatorname{Var}(Y^{x}).$
Next we show that for a fixed $\varepsilon>0$, we can always find a $x\in\mathbb{R}$ such that $\operatorname{Var}(Y^{x})\leq 2\varepsilon^{2}$. Since the support of $\mathbb{P}^{*}$ is $[0,1]$, we can bound
$\operatorname{Var}(Y^{x})=\mathbb{E}[(Y^{x}-\mathbb{E}[Y^{x}])^{2}]\leq\mathbb{E}[(Y^{x}-1)^{2}]=\mathbb{E}[(Y^{x}-1)^{2}\mathds{1}_{\{Y\geq 1-\varepsilon\}}]+\mathbb{E}[(Y^{x}-1)^{2}\mathds{1}_{\{Y<1-\varepsilon\}}]\leq\varepsilon^{2}+\mathbb{P}(Y^{x}<1-\varepsilon).$
Then
$\mathbb{P}(Y<1-\varepsilon)=\frac{\int_{0}^{1-\varepsilon}\exp(-\frac{(x-ty)^{2}}{2\sigma_{t}^{2}})\;\mathrm{d}y}{\int_{0}^{1}\exp(-\frac{(x-ty)^{2}}{2\sigma_{t}^{2}})\;\mathrm{d}y}\leq\frac{\int_{0}^{1-\varepsilon}\exp(-\frac{(x-t(1-\varepsilon))^{2}}{2\sigma_{t}^{2}})\;\mathrm{d}y}{\int_{1-\varepsilon/2}^{1}\exp(-\frac{(x-t(1-\varepsilon/2))^{2}}{2\sigma_{t}^{2}})\;\mathrm{d}y}=\frac{2(1-\varepsilon)}{\varepsilon}\exp\Big(-\frac{xt\varepsilon-t^{2}(\varepsilon-3\varepsilon^{2}/4)}{2\sigma_{t}^{2}}\Big).$
Thus, for every $\varepsilon$, there is a $x\in\mathbb{R}$ such that $\mathbb{P}(Y<1-\varepsilon)\leq\varepsilon^{2}$. Hence, for every $\varepsilon>0$, we obtain
$\partial_{x}v_{t}(x)=\frac{\sigma_{t}^{\prime}}{\sigma_{t}}+\Big(\frac{\sigma_{t}^{\prime}}{\sigma_{t}^{3}}t^{2}-\frac{t}{\sigma_{t}^{2}}\Big)O(\varepsilon^{2})\quad\Longrightarrow\quad\sup_{x\in\mathbb{R}}|\partial_{x}v_{t}(x)|\gtrsim\Big|\frac{\sigma_{t}^{\prime}}{\sigma_{t}}\Big|\quad\Longrightarrow\quad\Gamma_{t}\gtrsim\Big|\frac{\sigma_{t}^{\prime}}{\sigma_{t}}\Big|.\qed$
6.2 Proofs of Section˜4
Proof of Theorem˜4.2.
@@ 6.2 Proofs of Section˜4
Then we use Lemma˜A.1,
$\mathbb{E}_{\begin{subarray}{c}t\sim\mathcal{U}[0,1]\\ X_{t}\sim p_{t}\end{subarray}}\big[|v_{t}(X_{t})-\hat{v}_{t}(X_{t})|^{2}\big]=\mathbb{E}_{\begin{subarray}{c}t\sim\mathcal{U}[0,1],\\ Y\sim p^{*},\\ X_{t}\sim p_{t}\left(\cdot|Y\right)\end{subarray}}\big[\left|\hat{v}_{t}(X_{t})-v_{t}\left(X_{t}|Y\right)\right|^{2}\big]-C,\;\text{with}\;C\coloneqq\mathbb{E}_{\begin{subarray}{c}t\sim\mathcal{U}[0,1],\\ Y\sim p^{*},\\ X_{t}\sim p_{t}\left(\cdot|Y\right)\end{subarray}}[|v_{t}(X_{t})-v_{t}(X_{t}|Y)|^{2}].$ (33)
@@
$\mathbb{E}_{\begin{subarray}{c}t\sim\mathcal{U}[0,1],\\ Y\sim p^{*},\\ X_{t}\sim p_{t}\left(\cdot|Y\right)\end{subarray}}\big[\left|\hat{v}_{t}(X_{t})-v_{t}\left(X_{t}|Y\right)\right|^{2}\big]=\mathbb{E}_{Y\sim\mathbb{P}^{*}}[g(\hat{v},Y)\mathds{1}_{Y\in A}]+\mathbb{E}_{Y\sim\mathbb{P}^{*}}[g(\hat{v},Y)\mathds{1}_{Y\in\mathbb{R}^{d}\setminus A}]$
@@
$|v_{t}^{j}(x)|$ $=\Big|\int v^{j}_{t}(x|y)\frac{p_{t}(x|y)}{\int p_{t}(x|z)p^{*}(z)\;\mathrm{d}z}p^{*}(y)\;\mathrm{d}y\Big|=\Big|\int\Big(\frac{\sigma^{\prime}_{t}}{\sigma_{t}}x_{j}+\Big(1-\frac{\sigma^{\prime}_{t}}{\sigma_{t}}t\Big)y_{j}\Big)\frac{p_{t}(x|y)}{\int p_{t}(x|z)p^{*}(z)\;\mathrm{d}z}p^{*}(y)\;\mathrm{d}y\Big|$
@@
$\leq\Big(\Big|\frac{\sigma^{\prime}_{t}}{\sigma_{t}}\Big|+e^{2L}\Big|\Big(1-\frac{\sigma^{\prime}_{t}}{\sigma_{t}}t\Big)\Big|\frac{t}{t^{2}+\sigma_{t}^{2}}\Big)|x_{j}|+\Big|\Big(1-\frac{\sigma^{\prime}_{t}}{\sigma_{t}}t\Big)\Big|\Big(1+\frac{t^{2}}{\sigma_{t}^{2}}\Big)^{-\frac{1}{2}}.$
As $\max_{t\in[0,1]}\big|\frac{\sigma^{\prime}_{t}}{\sigma_{t}}\big|=\log(\sigma_{\min}^{-1})$, $(1+\frac{t^{2}}{\sigma_{t}^{2}})^{-\frac{1}{2}}\leq 1$ and due to Lemma˜A.2 $\frac{t}{t^{2}+\sigma_{t}^{2}}\leq\max(\log(\sigma_{\min}^{-1}),e^{2})$ we obtain for the maximum over $t\in[0,1]$ for $n$ big enough $|v_{t}^{j}(x)|\lesssim\log(n)^{2}|x_{j}|+\log(n).$ The constants lead to the bound in (15). An analogous calculation shows that, for $t\in[0,1]$, the norm of $v_{t}$ is bounded by
@@
$g(v,y)=\int_{0}^{1}\int|v_{t}(x)-v_{t}(x|y)|^{2}p_{t}(x|y)\;\mathrm{d}x\;\mathrm{d}t\lesssim\log(n)^{6}+\int_{0}^{1}\int|v(x|y)|^{2}p_{t}(x|y)\;\mathrm{d}x\;\mathrm{d}t.$
@@
$\int|v(x|y)|^{2}p_{t}(x|y)\;\mathrm{d}x=\int\Big|\frac{\sigma^{\prime}_{t}}{\sigma_{t}}x+\Big(1-\frac{\sigma^{\prime}_{t}}{\sigma_{t}}t\Big)y\Big|^{2}p_{t}(x|y)\;\mathrm{d}x=\Big|\frac{\sigma^{\prime}_{t}}{\sigma_{t}}\Big|^{2}\int|x|^{2}p_{t}(x|y)\;\mathrm{d}x+\Big(1-\frac{\sigma^{\prime}_{t}}{\sigma_{t}}t\Big)^{2}|y|^{2},$
and as $p_{t}(\cdot|y)$ is the density of a Gaussian with mean $y$ and variance $\sigma_{t}^{2}I_{d}$, which implies $\int|x|^{2}p_{t}(x|y)\;\mathrm{d}x=t^{2}|y|^{2}+\sigma_{t}^{2}d,$ we obtain for $g(v,y)$
$g(v,y)\lesssim\log(n)^{6}+\int_{0}^{1}\Big|\frac{\sigma^{\prime}_{t}}{\sigma_{t}}\Big|^{2}(t^{2}|y|^{2}+\sigma_{t}^{2}d)\;\mathrm{d}t+|y|^{2}\int_{0}^{1}\Big(1-\frac{\sigma^{\prime}_{t}}{\sigma_{t}}t\Big)^{2}\;\mathrm{d}t\lesssim\log(n)^{6}.$